# Mathematical Anatomy of Neutrino Decoherence in Red Turbulence: A Fractional Calculus Approach

### v1

@@ II.3 Relation to Fractional Calculus
The memory integral with power-law kernel exhibits a natural connection to fractional calculus for the physically most relevant case of red noise ($\nu<1$).
@@
For the Riemann-Liouville fractional integral of order $\alpha>0$ [6],
@@
Thus, for red noise spectra ($\nu<1$), the non-Markovian master equation incorporates a Riemann-Liouville fractional integral operator of order $1-\nu$. This fractional structure explains the emergence of Mittag-Leffler functions in the solution, which interpolate between initial power-law decay and eventual exponential relaxation.
Regularization for the general case.
For completeness, we note that the small-scale regularization $\exp(-\epsilon/u)$ ensures mathematical consistency for any $\nu$. However, in realistic turbulent environments, spectral indices are typically in the range $0<\nu\leq 1$, corresponding to red noise spectra where low-frequency fluctuations dominate. The case $\nu>1$ (blue noise) is physically less relevant for turbulence in supernovae, as it would imply energy concentrating at high frequencies, contrary to the observed inertial-range scaling of turbulent cascades.
@@ III Analytic Solution via Fractional Calculus
For $\nu<0$, the regularized correlation function exhibits a distinctive mathematical structure:
$K(t)=\xi\,t^{|\nu|}\exp\left(-\frac{\epsilon}{t}\right),\quad|\nu|=-\nu>0,$ (38)
where we have used the same regularization parameter $\epsilon$ as in Eq. (9) for consistency.
The Laplace transform yields:
$\tilde{K}_{\nu}(z)=\xi\,\Gamma(1-\nu)\,(z+\gamma)^{\nu-1},\quad\gamma=\frac{1}{\tau_{c}}.$ (39)
The key insight: The memory integral in the master equation becomes a fractional integral operator of order $|\nu|+1$.
From the definition of Riemann-Liouville fractional integral in Eq.35, our Master equation can be elegantly written as [6]:
$\frac{d}{dt}\rho(t)=-i[H_{0},\rho(t)]-\xi\Gamma(1-\nu)\,I^{|\nu|+1}[\sigma_{z},[\sigma_{z},\rho]].$ (40)
For $\nu=-1$ ($|\nu|=1$), this is a regular double integral. For $\nu=-2$ ($|\nu|=2$), it’s a triple integral.
III.1 Exact Solution via Operator Method
Consider the projection onto flavor coherence. Define $Q(t)=\langle\nu_{e}|\rho(t)|\nu_{x}\rangle$ in the matter basis. The equation simplifies to:
$\frac{d}{dt}Q(t)$ $=i\Delta_{m}Q(t)-4\xi\Gamma(1-\nu)I^{|\nu|+1}[Q](t),$ (41)
$Q(0)$ $=\frac{1}{2}\sin 2\theta_{m}\,.$ (42)
Applying the Laplace transform, we obtain
$(s-i\Delta_{m})\tilde{Q}(s)-Q(0)=-4\xi\Gamma(1-\nu)\frac{\tilde{Q}(s)}{s^{|\nu|+1}}\,.$ (43)
Thus:
$\tilde{Q}(s)=\frac{Q(0)}{s-i\Delta_{m}+4\xi\Gamma(1-\nu)s^{-(|\nu|+1)}}.$ (44)
III.2 Analytic Structure for Integer $|\nu|$
For $|\nu|=m\in\mathbb{Z}^{+}$, multiply both sides by $s^{m+1}$:
$\big[s^{m+2}-i\Delta_{m}s^{m+1}+4\xi\,m!\big]\tilde{Q}(s)\;=\;s^{m+1}Q(0).$ (45)
Inverse transforming yields an $(m+2)$-th order differential equation:
$Q^{(m+2)}(t)-i\Delta_{m}Q^{(m+1)}(t)-4\xi\Gamma(1+m)Q(t)=0.$ (46)
This is the key simplification: The nonlocal memory integral transforms into a local differential operator!
III.3 Mittag-Leffler Function Representation
The exact solution of the fractional master equation (40) can be expressed in terms of multivariate Mittag-Leffler functions. For the off-diagonal element $Q(t)$ in the high-density limit, one finds
$Q(t)=Q(0)\,E_{(|\nu|+2,\;|\nu|+1)}\Bigl(i\Delta_{m}t^{|\nu|+2},\;-4\xi\,\Gamma(1+|\nu|)\,t^{|\nu|+2}\Bigr),$ (47)
where the $(p,q)$-variate Mittag-Leffler function is defined by the double series
$E_{(p,q)}(a,b)\;=\;\sum_{n=0}^{\infty}\frac{a^{n}t^{pn}}{(pn)!}\;{}_{1}F_{1}\!\left(\begin{array}[]{c}1\\ qn+1\end{array};bt^{q}\right),$ (48)
with $_{1}F_{1}$ being the confluent hypergeometric function. This representation reveals the interplay between oscillatory dynamics (governed by the parameter $a=i\Delta_{m}$) and algebraic damping (governed by $b=-4\xi\,\Gamma(1+|\nu|)$).
III.4 Physical Interpretation: Accelerated Decoherence
The fractional integral operator $I^{|\nu|+1}$ represents a smearing of quantum phase information over an algebraically weighted past. This leads to:
@@ III.4 Physical Interpretation: Accelerated Decoherence
    Enhanced memory effects: The kernel $t^{|\nu|}$ gives greater weight to recent history compared to white noise.
@@
    Algebraic-to-exponential crossover: At $t\sim\tau_{c}$, the exponential cutoff $e^{-t/\tau_{c}}$ dominates, recovering Markovian behavior.
@@
    Spectral reddening: The Fourier transform $\tilde{K}_{\nu}(\omega)\sim|\omega|^{-(|\nu|+1)}$ shows enhanced low-frequency power.
@@ IV Conclusion
The central achievement of this work is the development of a physically motivated regularization scheme for the ultraviolet divergence that arises in power-law correlation functions when the spectral index $\nu\geq 1$. By introducing an exponential cutoff $\exp(-\epsilon/t)$ at the dissipation time scale $\epsilon$, we ensure mathematical consistency of all integrals and operators while preserving the essential power-law behavior in the inertial range. This regularization enables a well-defined fractional calculus representation of the exact non-Markovian master equation, obtained through the Nakajima-Zwanzig projection technique for classical Gaussian noise.
@@
Our analysis elucidates the profound connection between neutrino flavor evolution and fractional calculus. For red turbulence spectra ($\nu<1$), the memory kernel corresponds to a Riemann-Liouville fractional integral of order $1-\nu$, while for extremely red spectra ($\nu<0$) it becomes a higher-order fractional operator. This fractional structure explains the characteristic non-exponential decoherence, with an initial power-law decay $P_{ee}(x)\sim 1-C_{\nu}x^{2-\nu}$ transitioning to exponential damping at late times.
We provided the first detailed study of neutrino decoherence for negative spectral exponents ($\nu<0$), corresponding to extremely red turbulence spectra. In these cases, the memory integral translates into higher-order fractional operators, leading to distinct dynamical scaling and mathematical features. The exact solutions for these regimes demonstrate how spectral reddening affects decoherence efficiency and memory effects.
The physical implications for supernova neutrinos are significant. Our exact analytical results serve as a crucial benchmark for numerical simulations and provide clear guidelines: (i) the high-density matter basis should be used when $\Delta_{m}\gg\Delta_{v}$ to properly account for matter effects; (ii) the small-scale cutoff must be included for spectral indices $\nu\geq 1$ to ensure mathematical rigor; (iii) the fractional calculus formulation offers efficient computational methods through established properties of Mittag-Leffler functions; and (iv) the interplay between spectral index, regularization scale, and decoherence efficiency can now be systematically studied within our exact framework.
@@
In summary, we have presented a complete exact mathematical framework for analyzing neutrino decoherence in power-law correlated turbulent environments. By combining rigorous regularization techniques with fractional calculus methods, we have developed closed-form analytical solutions that capture the full interplay between coherent oscillations and algebraic decoherence induced by long-range temporal correlations. Our work establishes firm mathematical foundations for future studies of neutrino-turbulence interactions in astrophysical environments, particularly in the complex conditions of core-collapse supernovae. The fractional calculus perspective not only provides practical computational tools but also reveals deep connections between neutrino flavor evolution and other physical systems exhibiting anomalous dynamics. As supernova simulations continue to increase in sophistication and next-generation neutrino detectors come online, the exact analytical treatment of turbulence effects developed here will prove essential for accurate interpretation of neutrino signals as probes of extreme astrophysical phenomena.
Acknowledgements.
We thank Thierry Foglizzo for insightful discussions and the developers of the mpmath Python library for high-precision numerical evaluation of special functions. During the preparation of this work, the authors used DeepSeekV3 to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication. This work is supported by the Fundamental Research Funds for the Central Universities under No. 020114380057, and by K. C. Wong Educational Foundation. A.A. is supported by the National Science Foundation of China (NSFC) through the grant No. 12350410358; the Talent Scientific Research Program of College of Physics, Sichuan University, Grant No. 1082204112427; the Fostering Program in Disciplines Possessing Novel Features for Natural Science of Sichuan University, Grant No.2020SCUNL209 and 1000 Talent program of Sichuan province 2021.

### v3

@@ II.3 Relation to Fractional Calculus
The memory integral with power-law kernel exhibits a natural connection to fractional calculus throughout the long-memory regime $\nu<1$. In our convention, red noise corresponds to $\nu<0$, while $\nu=1$ is the white-noise boundary.
@@
For the Riemann-Liouville fractional integral of order $\alpha>0$ Balachandran (2023),
@@
Thus, for $\nu<1$, the non-Markovian master equation incorporates a Riemann-Liouville fractional integral operator of order $1-\nu$. At $\nu=1$ (white-noise limit), this order tends to zero and the dynamics approaches the local Markovian form. For red spectra ($\nu<0$), the order exceeds unity and memory effects are enhanced. This fractional structure explains the emergence of Mittag-Leffler functions in the solution.
Auxiliary Regularization and Renormalized Interpretation.
For completeness, we note that the auxiliary short-time regularization $\exp(-\epsilon/u)$ ensures mathematical consistency of the intermediate kernel representation for any $\nu$. In the renormalized treatment developed in Sec. III, ultraviolet local terms are absorbed into $\Delta_{m}^{\mathrm{ren}}$. In our convention, red spectra correspond to $\nu<0$, $\nu=1$ is the white-noise boundary, and $\nu>1$ corresponds to blue noise.
@@ III Analytic Solution via Fractional Calculus
The exact construction in Sec. II used an explicit short-time regulator to make every intermediate integral convergent. For analytic continuation in $\nu$, however, a pure exponential truncation mixes scales and obscures the fractional structure of the kernel. In this section we instead adopt a renormalization scheme: the ultraviolet local piece is absorbed into a frequency shift, while the nonlocal power-law part keeps the Mittag-Leffler form of the solution.
To remain consistent with Sec. II, we keep the same effective turbulence parameter $\kappa_{\nu}$ from Eq. (19).
III.1 Hadamard Finite-Part Kernel
We define the memory kernel as the finite-part distribution
$t_{+}^{-\nu}\equiv\operatorname{Fp}\!\left[\frac{t^{-\nu}\Theta(t)}{\Gamma(1-\nu)}\right],$ (39)
and write
$K(t)=\xi_{\nu}\,t_{+}^{-\nu}.$ (40)
Equation (39) means that the singular $t\to 0^{+}$ part is treated in the Hadamard finite-part sense, i.e. only the finite distributional contribution is retained after subtracting the divergent local terms on test functions.
Its Laplace transform is obtained by analytic continuation:
$\tilde{K}_{\nu}(s)=\xi_{\nu}\,s^{\nu-1},\qquad(\text{equivalently }\tilde{K}_{\nu}(s)=\xi_{\nu}\Gamma(1-\nu)s^{\nu-1}\text{ before normalization}).$ (41)
III.2 Renormalized Laplace-Space Equation
For the off-diagonal coherence amplitude $Q(t)$, the Laplace equation is
$(s-i\Delta_{m})\tilde{Q}(s)-Q(0)=-\kappa_{\nu}s^{\nu-1}\tilde{Q}(s).$ (42)
For $\nu>1$, $s^{\nu-1}$ contains an ultraviolet local piece. Introduce a renormalization scale $\mu$:
$s^{\nu-1}=\bigl[s^{\nu-1}-\mu^{\nu-1}\bigr]+\mu^{\nu-1},$ (43)
and absorb the constant term into the oscillation frequency:
$\Delta_{m}^{\mathrm{ren}}(\mu)\equiv\Delta_{m}+i\kappa_{\nu}\mu^{\nu-1}.$ (44)
Taking the finite renormalized limit defines $\Delta_{m}^{\mathrm{ren}}$, and Eq. (42) becomes
$\boxed{\tilde{Q}(s)=\frac{Q(0)}{s-i\Delta_{m}^{\mathrm{ren}}+\kappa_{\nu}s^{\nu-1}}}$ (45)
which preserves the pure fractional power kernel.
III.3 Integer Branch and Local Representation
For $\nu=-m$ with $m\in\mathbb{Z}^{+}$, the analytically continued form yields
$\big[s^{m+2}-i\Delta_{m}^{\mathrm{ren}}s^{m+1}+\kappa_{\nu}\,m!\big]\tilde{Q}(s)\;=\;s^{m+1}Q(0).$ (46)
Inverse Laplace transformation gives the equivalent local equation
$Q^{(m+2)}(t)-i\Delta_{m}^{\mathrm{ren}}Q^{(m+1)}(t)-\kappa_{\nu}\Gamma(1+m)Q(t)=0.$ (47)
III.4 Mittag-Leffler Solution
The renormalized fractional equation keeps the standard Mittag-Leffler structure:
$\boxed{Q(t)=Q(0)\,E_{\nu,1}\!\left(i\Delta_{m}^{\mathrm{ren}}t^{\nu}-\kappa_{\nu}t^{\nu}\right)}$ (48)
with
$E_{\nu,1}(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{\Gamma(\nu n+1)}.$ (49)
Thus the non-exponential relaxation law is retained exactly after renormalization.
III.5 Physical Meaning of Renormalization
The renormalized solution has a direct physical interpretation:
@@ III.5 Physical Meaning of Renormalization
    $\Delta_{m}^{\mathrm{ren}}$ contains the turbulence-induced level shift (Lamb-shift-like contribution from ultraviolet local terms).
@@
    $\kappa_{\nu}$ controls genuine nonlocal decoherence, giving Mittag-Leffler rather than purely exponential damping.
@@
    The same mathematical structure appears in anomalous diffusion with memory kernels, establishing a one-to-one mapping between neutrino decoherence and fractional transport dynamics.
III.6 Outlook: Inclusion of Neutrino-Neutrino Interactions
The present analysis has focused on neutrino decoherence induced by turbulent density fluctuations, treating neutrinos as test particles propagating through a stochastic medium. In core-collapse supernovae, however, the high neutrino densities give rise to additional collective effects through neutrino-neutrino interactions Duan et al. (2010b, a). These interactions introduce nonlinearity into the evolution equations and can trigger fast flavor instabilities on small scales Bhattacharyya and Dasgupta (2021); Nagakura (2023); Ehring et al. (2023b), with recent works identifying new instability mechanisms Fiorillo et al. (2025) and revealing the quantum structure of collective flavor conversions Fiorillo and Raffelt (2025). The occurrence and impact of these instabilities in realistic supernova environments have been extensively studied in multidimensional simulations Abbar et al. (2019), and comprehensive reviews now synthesize our understanding of neutrino oscillations in both core-collapse supernovae and neutron star mergers Johns et al. (2025). These developments underscore the need to understand how turbulence and self-interactions jointly shape flavor evolution.
In the presence of both turbulence and neutrino self-interactions, the master equation acquires additional structure. The total Hamiltonian becomes $H_{\mathrm{total}}=H_{0}+H_{\nu\nu}(\rho)+V(t)\sigma_{z}$, where $H_{\nu\nu}(\rho)$ depends nonlinearly on the neutrino density matrix itself. Applying the Nakajima-Zwanzig projection technique to this case yields a more complex integro-differential equation where the memory kernel now depends on the history of $\rho$ through both the turbulent correlator and the self-interaction term. For power-law correlated turbulence, the resulting equation combines fractional memory with nonlinear coupling, leading to phenomena such as turbulence-enhanced or turbulence-suppressed collective oscillations.
A detailed treatment of this coupled system, including the interplay between fractional decoherence and collective flavor evolution, is developed in our companion paper Bao and Addazi (2026). The fractional calculus framework established here provides the mathematical foundation for that analysis, demonstrating that the Mittag-Leffler structure persists even in the nonlinear regime under appropriate approximations.
@@ IV Conclusion
The central achievement of this work is the development of a physically motivated renormalization scheme for the ultraviolet divergence that arises in power-law correlation functions when the spectral index $\nu\geq 1$. Using the Hadamard finite-part definition of the kernel and analytic continuation in Laplace space, we isolate the ultraviolet local term and absorb it into a renormalized oscillation frequency $\Delta_{m}^{\mathrm{ren}}$. This procedure preserves the pure fractional power-law kernel and therefore the Mittag-Leffler structure of the exact non-Markovian solution obtained from the Nakajima-Zwanzig framework.
@@
Our analysis elucidates the profound connection between neutrino flavor evolution and fractional calculus. For red turbulence spectra ($\nu<0$), the memory kernel corresponds to a higher-order fractional operator, while $\nu=1$ corresponds to the white-noise boundary. More generally, $\nu<1$ yields a Riemann-Liouville fractional memory structure of order $1-\nu$. This fractional structure explains characteristic non-exponential decoherence and long-memory relaxation described by Mittag-Leffler functions.
We provided the first detailed study of neutrino decoherence for negative spectral exponents ($\nu<0$), corresponding to red turbulence spectra. In these cases, the memory integral translates into higher-order fractional operators, leading to distinct dynamical scaling and mathematical features. The exact solutions for these regimes demonstrate how spectral reddening affects decoherence efficiency and memory effects.
The physical implications for supernova neutrinos are significant. Our exact analytical results serve as a crucial benchmark for numerical simulations and provide clear guidelines: (i) the high-density matter basis should be used when $\Delta_{m}\gg\Delta_{v}$ to properly account for matter effects; (ii) for spectral indices $\nu\geq 1$, ultraviolet local terms should be treated by renormalization into $\Delta_{m}^{\mathrm{ren}}$ rather than by inserting an external exponential truncation; (iii) the fractional calculus formulation offers efficient computational methods through established properties of Mittag-Leffler functions; and (iv) the interplay between spectral index, renormalization scale, and decoherence efficiency can now be systematically studied within our exact framework.
@@
In summary, we have presented a complete exact mathematical framework for analyzing neutrino decoherence in power-law correlated turbulent environments. By combining rigorous renormalization techniques with fractional calculus methods, we have developed closed-form analytical solutions that capture the full interplay between coherent oscillations and algebraic decoherence induced by long-range temporal correlations. Our work establishes firm mathematical foundations for future studies of neutrino-turbulence interactions in astrophysical environments, particularly in the complex conditions of core-collapse supernovae. The fractional calculus perspective not only provides practical computational tools but also reveals deep connections between neutrino flavor evolution and other physical systems exhibiting anomalous dynamics. As supernova simulations continue to increase in sophistication and next-generation neutrino detectors come online, the exact analytical treatment of turbulence effects developed here will prove essential for accurate interpretation of neutrino signals as probes of extreme astrophysical phenomena.
We thank Thierry Foglizzo for insightful discussions and the developers of the mpmath Python library for high-precision numerical evaluation of special functions. During the preparation of this work, the authors used DeepSeekV3 to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication. This work is supported by the Fundamental Research Funds for the Central Universities under No. 020114380057, and by K. C. Wong Educational Foundation. $A.$ is supported by the National Science Foundation of China (NSFC) through the grant No. 12350410358; the Talent Scientific Research Program of College of Physics, Sichuan University, Grant No. 1082204112427; the Fostering Program in Disciplines Possessing Novel Features for Natural Science of Sichuan University, Grant No.2020SCUNL209 and 1000 Talent program of Sichuan province 2021. S.Z. is supported by the National Science Foundation of China (NSFC) through the grant Nos. 12288102, 12473031, 12393811; the Yunnan Revitalization Talent Support Program–Young Talent project; and the Yunnan Fundamental Research Project No. 202501AS070078.
{contribution}
All authors contributed equally.